OpenStudy (anonymous):

a watch loses 3 minutes ever 24 hours. how much time will it lose in 2 hours. ( please walk me through this step by step)

OpenStudy (wolf1728):

If it loses 3 minutes every 24 hours then it loses (3/24) minutes every hour. = 1/8 minute per hour = (1/8) * 60 = 7.5 seconds per hour So in two hours it loses 15 seconds.

OpenStudy (anonymous):

Thanks your a life saver!

OpenStudy (wolf1728):

Glad to help out

OpenStudy (anonymous):

you're*